OpenStudy (anonymous):

please, help me with this solve each exponential equation s:1.) e^x-2 = e^3x+1 2.) 7 = 6 + e^3-2x

OpenStudy (anonymous):

\[e\large^x-2=e^{3x}+1\] or \[\large e^{x-2}=e^{3x}+1\]?

hartnn (hartnn):

lol, i guess e^(x-2) = e^(3x+1)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

in that case i can do it

OpenStudy (anonymous):

please. thankyou so much

OpenStudy (anonymous):

which one is it? there is some ambiguity here

hartnn (hartnn):

so, we just compare the exponents of 'e' e^(x-2) = e^(3x+1) so, x-2 = 3x+1 can you find x from here ?

hartnn (hartnn):

lets assume the easiest one ;)

OpenStudy (anonymous):

yes, i end up solving there.

OpenStudy (anonymous):

is the second one \[\large 7=6+e^{3-2x}\]?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if so, subtract \(6\) and get \[e^{3-2x}=1\] making \[3-2x=0\]solve for \(x\)

OpenStudy (anonymous):

oh, okayy thanks! i got it

hartnn (hartnn):

since you are new here \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

Thankyou! :)

OpenStudy (anonymous):

why 3 - 2x = 0?

hartnn (hartnn):

\(e^0 = 1 \\ e^{3-2x}=1 \\ 3-2x = 0\)

hartnn (hartnn):

infact (anything except 0)^0 =1

OpenStudy (anonymous):

Ah, okayy Thanks! how about this one? (5/2)^x = 4/25

OpenStudy (anonymous):

please, if it's okay

OpenStudy (anonymous):

i still find it hard

hartnn (hartnn):

hey, no problem at all :)

hartnn (hartnn):

4/25 = ...? can you express 4 and 25 as powers of 2 and 5 ?

OpenStudy (anonymous):

2^2

OpenStudy (anonymous):

5^5

hartnn (hartnn):

so, 4/25 = (2/5)^2 agree ?

hartnn (hartnn):

5^5 ? O.o

OpenStudy (anonymous):

ayy, joke haha/ it's 5^2!

OpenStudy (anonymous):

^_^

hartnn (hartnn):

yes, so (2/5)^2 but we have (5/2)^x on left so,

hartnn (hartnn):

\(\Huge (\dfrac{2}{5})^2 = \dfrac{1}{(\dfrac{5}{2})^2}= (\dfrac{5}{2})^{-2}\) got this ?

hartnn (hartnn):

there ?

OpenStudy (anonymous):

are you already a professor?

OpenStudy (anonymous):

Thanks. gonna take note with this

hartnn (hartnn):

i am a student just like you ^_^

hartnn (hartnn):

and could you find x there ?

OpenStudy (anonymous):

what year are you in?

OpenStudy (anonymous):

wait a minute

hartnn (hartnn):

i already completed my studies :)

OpenStudy (anonymous):

so ,the x on the first problem is -3/2? and the x on my second problem would be -2. ? is my answer correct?

OpenStudy (anonymous):

heyy heyy

hartnn (hartnn):

for 2nd problem. 3-2x = 0 right ? x = 3/2

hartnn (hartnn):

ok

OpenStudy (anonymous):

for my question e^(x-2) = e^(3x+1) we end up at x -2 = 3x+1 , can u explain to me how we come up the answer of 1.5?

hartnn (hartnn):

its -1.5 x -2 = 3x+1 -2 = 2x +1 -2-1 = 2x 2x = -3 x=-3/2 = -1.5

OpenStudy (anonymous):

ah, okay. really big thanks! is it really okay, i still have more questions

OpenStudy (anonymous):

:)

hartnn (hartnn):

i'll be glad to help :)